OpenStudy (anonymous):

A small amount of triglyceride, 24 mg, decomposes in a first order reaction to 12 mg in 20 minutes. What is the value of the first order rate constant?

OpenStudy (aaronq):

use the first order exponential decay formula.

OpenStudy (anonymous):

I don't think I've learned that yet. I'm trying to use the first order integrated rate law but I keep coming up with the wrong answer and I'm not sure what I'm doing wrong

OpenStudy (aaronq):

you can use this formula \([A]_t=[A_o]*e^{kt}\) \(12=24*e^{k*20~min} \rightarrow k= \dfrac{20~min}{ln\dfrac{12}{24}}\)

OpenStudy (anonymous):

I calculated that and got a negative answer which doesn't match any of the option. I also tried \[\ln[24]=-k(20)+\ln[12]\]but that didn't work either and I've been stuck there

OpenStudy (aaronq):

i made a mistake, it should be \(k=-\dfrac{ln\dfrac{12}{24}}{20~min}=0.03465735~/min\) when you use the integrated rate law the "intercept" in the original amount, so your equation should read: ln[12]=-k(20 min) + ln[24] which gives you the same thing as i wrote.